OpenStudy (anonymous):

Solve the system by the elimination method. 6x-9y=7 -4x+6y=10 a. An infinite number of solutions b. (0, 5) c. (5/6,5/2) d. No solution e. None of the above

OpenStudy (helder_edwin):

try to have the coeficients of x or y equal

OpenStudy (helder_edwin):

for example multiply the first eq by 4 and the second by 6 what do you get?

OpenStudy (anonymous):

Oh, OK thanks

OpenStudy (theviper):

@Kalisa @helder_edwin d.No Solution as\[\frac{a_1}{a_2}=\frac{b_1}{b_2}\neq \frac{c_1}{c_2}\]

OpenStudy (anonymous):

d. No Solution\[\left\{y=\frac{2 x}{3}-\frac{7}{9},y=\frac{2 x}{3}+\frac{5}{3}\right\} \]Both lines have the same slope, 2/3, and will never cross when plotted.

OpenStudy (anonymous):

@TheViper Thank you for the medal.

OpenStudy (theviper):

@robtobey i m ur fan also yw:)

OpenStudy (anonymous):

Thank you again TheViper

OpenStudy (theviper):

yw:)